How do you solve log little two underneath 8X^3
Answer by Edwin McCravy(20086)

log2(8x³) =
log2(2³x³) =
log2(2³) + log2(x³) =
3·log2(2) + 3·log2(x) =
3·1 + 3·log2(x) =
3 + 3·log2(x) =
3[1 + log2(x)]

Edwin